If the quantity of avocados demanded changes from 2000 boxes to 3000 boxes per week because the price has changed from R700 to R500, the price elasticity of demand for avocados will be...


1
Expert's answer
2021-06-22T16:57:14-0400

The price elasticity of demand for avocados will be:

"Ed = \frac{3000-2000} {500-700} ×\frac{500+700} {3000+2000} = -1.2,"

so the demand is elastic.